DOES WATER PREFER TO DONATE A PROTON TO AN F OR TO a Cl ATOM? - A ROTATIONAL STUDY OF CH_{3}CHClF...H_{2}O

GANG FENG, LUCA EVANGELISTI and \underline{W. CAMINATI}, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy; \underline{LAURA B. FAVERO}, Istituto per lo Studio dei Materiali Nanostrutturati (ISMN, Sezione di Bologna), CNR, Via Gobetti 101, I-40129 Bologna, Italy; \underline{JENS-UWE GRABOW}, Lehrgebiet Physikalische Chemie A, Institut für Physikalische Chemie und Elektrochemie, Universtät Hannover, Callinstr. 3A, D-30167 Hannover, Germany; ZHINING XIA, Chemistry and Chemistry Engineering College, Chongqing University, Chongqing, 400030, P. R. China.

We measured the molecular beam Fourier transform microwave spectra of six isotopologues of the 1:1 adduct of CH_{3}CHClF with water. The water prefers to form an O-H...F rather than an O-H...Cl hydrogen bond. This is exactly the contrary of what observed in the chlorofluoromethane-water adduct, where a O-H...Cl link was formed ^{a}. Besides the rotational constants, the quadrupole coupling constants of the chlorine atom have been determined. In addition, information on the internal dynamics has been obtained.

aW.Caminati, S.Melandri, A.Maris and P.Ottaviani, Angew. Chem. Int. Ed. 45, 2438-2442 (2006)